# Claim Detection for Automated Fact-checking: A Survey on Monolingual, Multilingual and Cross-Lingual Research

### 23 Jan 2024 v2

@@ Abstract
journal: Review
@@ 2.1 Definition of Claim
    Claim - A claim is defined as a statement containing a purported fact (Das et al., 2023). For example, the statement “team X is the best team in FIFA 2023” is an opinion, whereas, the statement “United States is one of the host countries of the next FIFA World Cup” is a claim that can be checked against an objective piece of evidence. Zeng et al. (2021) give a slightly different definition for a claim as a “factual statement under investigation”.
@@
    Check-worthy Claim - Verifying all the assertions about the real world is impractical, and hence it demands a prioritization process. This objective is generally handled as an estimation of check-worthiness of a verifiable claim in the literature. Due to the subjective nature of this task, it is hard to define the check-worthiness of a verifiable claim and it may vary according to the topic discussed in the claim and the user group who are interested in the claim. Further, the worthiness may vary over time (Guo et al., 2022), as the interest of a claim may fade or increase as stories develop, which makes it more challenging. Some of the common factors used to determine the check-worthiness of verifiable claims include the popularity of the claim, amount of public interest in the verdict of the claim, impact of the verdict, and timeliness of the verdict (Das et al., 2023; Micallef et al., 2022). Further claim types (e.g. attention-worthy claims, harmful claims) can be found in the literature depending on the criteria used to define the worthiness.
@@ 3.1 Datasets
Dataset Objective Label Language Topic Source Size
@@ 3.6 Discussion
Further, it can be observed that most of the studies on verifiable claim detection rely on pre-trained transformer models, especially BERT and its variations to obtain state-of-the-art performance with limited fine-tuning. Moreover, limited research has utilized tweet-specific features for the classification of Twitter data, and translation is used as a data augmentation technique to handle the limited training data issue. Further, it can be noticed that performing fine-grained verifiable claim-type identification has been carried out in the literature only in monolingual settings. This could be due to the unavailability of the training data in a multilingual environment and the challenges associated with developing multilingual data annotated with fine-grained claim types.
@@ 4.1 Datasets
Dataset Criteria Task Label Language Topic Source Size
@@ 4.5 Evaluation
Metrics such as accuracy, precision, recall, and F1-score (refer Section 3.5) are widely used as an evaluation measure when the prioritization is carried out as a classification task.
@@ 5.1 Datasets
Dataset Label Language Language Pair Topic Source Size
@@ 5.5 Evaluation
Evaluating claim matching depends on whether the problem is treated as a classification or ranking task. Evaluation metrics discussed in Section 3.5 for the classification tasks and in Section 4.5 for the ranking tasks can be applied according to the nature of the problem.
@@ 7 Open Challenges
    Limited Multilingual Datasets: One of the key aspects hindering the progress of multilingual claim detection research is the unavailability of training data. Especially, comprehensive multitopic claim detection datasets, verifiable claim type detection datasets, claim clustering datasets, and explainable claim detection are yet to be developed for the research progress even in monolingual settings.
@@ 8 Conclusion
As discussed previously, various factors affect the progress of multilingual claim detection research. Notably, the unavailability of datasets meeting the task requirement serves as a key challenge. Promising future direction includes the development of credible and comprehensive datasets, generalized solutions, explainable claim detection, and time-aware claim detection.

### 18 Mar 2024 v3

@@ Abstract
journal: Natrual Language Processing
@@ 2.1 Definition of Claim
(a) Black arrows in the figure indicate claims and red arrows indicate non-claims
@@
    Claim - A claim is defined as a statement containing a purported fact about the real world (Das et al., 2023). For example, the statement “team X is the best team in FIFA 2023” is an opinion, whereas, the statement “United States is one of the host countries of the next FIFA World Cup” is a claim that can be checked against an objective piece of evidence. Zeng et al. (2021) give a slightly different definition for a claim as a “factual statement under investigation”.
@@
    Check-worthy Claim - Verifying all the assertions about the real world is impractical, and hence it demands a prioritization process. This objective is generally handled as an estimation of check-worthiness of a verifiable claim in the literature. Due to the subjective nature of this task, it is hard to define the check-worthiness of a verifiable claim and it may vary according to the topic discussed in the claim and the user group who are interested in the claim. Further, the worthiness may vary over time (Guo et al., 2022), as the interest of a claim may fade or increase as stories develop, which makes it more challenging. Some of the common factors used to determine the check-worthiness of verifiable claims include the popularity of the claim, amount of public interest in the verdict of the claim, impact of the verdict, and timeliness of the verdict (Das et al., 2023; Micallef et al., 2022). In addition to the check-worthiness, the following criteria are used in the literature for prioritizing claims:
    • (a)
      Attention-worthy Claims: A verifiable claim that should get the attention of the policymakers and government entities (Nakov et al., 2022; Shaar et al., 2021a).
    • (b)
      Harmful Claims: A verifiable claim that is harmful to society (Nakov et al., 2022; Shaar et al., 2021a).
    • (c)
      Interesting to the general public: A verifiable claim that may have an impact on society or attract interest from the general public (Shaar et al., 2021a).
@@ 3.1 Datasets
Dataset Objective Label Language Topic Source Size Evaluation
@@ 3.5 Evaluation
Table 1 shows the evaluation metrics used in the existing verifiable claim detection datasets. It can be observed that precision, recall, and F1 score are widely used for the evaluation with the exception of the CheckThat 2022 dataset (Nakov et al., 2022). This is due to the fairly balanced nature of the CheckThat 2022 dataset, where the authors recommended reporting only accuracy as the evaluation metric.
@@ 3.6 Discussion
Further, it can be observed that most of the studies on verifiable claim detection rely on pre-trained transformer models, especially BERT and its variations to obtain state-of-the-art performance with limited fine-tuning. Moreover, limited research has utilized tweet-specific features for the classification of Twitter data, and translation is used as a data augmentation technique to handle the limited training data issue. Further, it can be noticed that performing fine-grained verifiable claim-type identification has been carried out in the literature only in monolingual settings. This could be due to the unavailability of the training data in a multilingual environment and the challenges associated with developing multilingual data annotated with fine-grained claim types. Moreover, very little effort has been made to date in the literature towards utilizing large language models (LLMs) for claim detection. We consider however that employing LLMs for claim detection represents an interesting avenue for future research.
@@ 4.1 Datasets
Dataset Criteria Task Label Language Topic Source Size Evaluation
@@ 4.5 Evaluation
Metrics such as accuracy, precision, recall, and F1-score (refer Section 3.5) are widely used as an evaluation measure when the prioritization is carried out as a classification task. Table 2 shows the evaluation metrics used in the existing claim prioritization datasets. It can be observed that a wide selection of metrics is used in the literature depending on whether the task is modeled as a classification or ranking problem. It is worth noting that the CheckThat 2022 dataset (Nakov et al., 2022) recommended reporting the F1 score to account for class imbalance.
@@ 4.6 Discussion
Table 4: Summary of models used for claim detection and prioritization research.
Model Type Model
Reference
Machine Learning Logistic Regression
Kartal (2020); Panda and Levitan (2021); Konstantinovskiy et al. (2021); Beltrán et al. (2021)
Random Forest
Smeros et al. (2021); Tarannum et al. (2022)
Support Vector Machine
Konstantinovskiy et al. (2021); Schlicht et al. (2021); Beltrán et al. (2021); Tarannum et al. (2022)
Deep Learning LSTM
Prabhakar et al. (2020)
Bi-LSTM
Martinez-Rico et al. (2020); Rony et al. (2020)
CNN
Martinez-Rico et al. (2020)
Feed Forward NN
Martinez-Rico et al. (2020); Konstantinovskiy et al. (2021)
Falir
Woloszyn et al. (2021)
Transformer Family BERT
Kartal (2020); Prabhakar et al. (2020); Cheema et al. (2020); Alam et al. (2020); Panda and Levitan (2021); Pathak and Srihari (2021); Zhou et al. (2021); Uyangodage et al. (2021); Smeros et al. (2021); Alam et al. (2021); Zengin et al. (2021); Hasanain and Elsayed (2022); Savchev (2022); Kartal and Kutlu (2022); Suri and Dudeja (2022); Eyuboglu et al. (2023); Sadouk et al. (2023)
mBERT
Hasanain and Elsayed (2020); Alam et al. (2020); Panda and Levitan (2021); Uyangodage et al. (2021); Alam et al. (2021); Zengin et al. (2021); Hasanain and Elsayed (2022); Kartal and Kutlu (2022); Tarannum et al. (2022); Schlicht et al. (2023); Sadouk et al. (2023)
XLM-r
Alam et al. (2020, 2021); Beltrán et al. (2021); Hüsünbeyi et al. (2022); Tarannum et al. (2022); Du et al. (2022); Schlicht et al. (2023); Aziz et al. (2023)
RoBERTa
Alam et al. (2020); Zhou et al. (2021); Alam et al. (2021); Woloszyn et al. (2021); Williams et al. (2021); Savchev (2022); Sadouk et al. (2023); Eyuboglu et al. (2023)
DistilBERT
Prabhakar et al. (2020); Savchev (2022); Eyuboglu et al. (2023)
BERTweet
Zhou et al. (2021); Woloszyn et al. (2021); Hüsünbeyi et al. (2022); Aziz et al. (2023)
SentenceBERT
Schlicht et al. (2021)
ConvBERT
Hüsünbeyi et al. (2022)
ALBERT
Alam et al. (2021); Eyuboglu et al. (2023); Sadouk et al. (2023)
XL-Net
Kartal and Kutlu (2022); Sadouk et al. (2023)
Transformer
Panda and Levitan (2021); Du et al. (2022); Gollapalli et al. (2023)
FastText
Kartal (2020); Alam et al. (2021)
AraBERT
Alam et al. (2020); Hussein et al. (2021); Henia et al. (2021); Zengin et al. (2021); Hasanain and Elsayed (2022); Aziz et al. (2023); Eyuboglu et al. (2023)
Spanish BERT
Uyangodage et al. (2021)
BERTurk
Uyangodage et al. (2021); Zengin et al. (2021); Hasanain and Elsayed (2022); Hüsünbeyi et al. (2022); Eyuboglu et al. (2023)
DutchBERT
Alam et al. (2020)
SalvicBERT
Hasanain and Elsayed (2022)
Large Language Models GPT
Agrestia et al. (2022); Sadouk et al. (2023)
Table 4 summarizes the models used in the literature for claim detection and claim prioritization. It can be noticed that transformer-based models are widely used compared to machine learning approaches and other deep learning learning approaches. Especially the BERT model and its architectural variations such as RoBERTa, XLM-r, ALBERT, BERTweet, Sentence BERT, DistilBERT, and ConvBERT, its multilingual variations such as mBERT, and the language-specific variations such as AraBERT, Spanish BERT, and BERTurk are commonly applied. Further, it can be noticed that large language models are yet to be explored in this line of research.
@@ 5.1 Datasets
Dataset Label Language Language Pair Topic Source Size Evaluation
@@ 5.5 Evaluation
Evaluating claim matching depends on whether the problem is treated as a classification or ranking task. Evaluation metrics discussed in Section 3.5 for the classification tasks and in Section 4.5 for the ranking tasks can be applied according to the nature of the problem. Table 5 shows the evaluation metrics used in the existing claim-matching datasets. It can be observed that Mean Reciprocal Ranking (MRR) and Mean Average Precision@K (MAP@K) are widely used for evaluation. Interestingly MultiClaim (Pikuliak et al., 2023) reports Precision@K and Recall@K as the ranking measures.
@@ 7 Open Challenges
    Limited Multilingual Datasets: One of the key aspects hindering the progress of multilingual claim detection research is the unavailability of training data. Especially, comprehensive multitopic claim detection datasets, verifiable claim type detection datasets, claim clustering datasets, and explainable claim detection are yet to be developed for the research progress even in monolingual settings. Automated generative approaches may be used as an alternative to generating claim detection datasets (Bussotti et al., 2023; Veltri et al., 2023).
@@ 8 Conclusion
As discussed previously, various factors affect the progress of multilingual claim detection research. Notably, the unavailability of datasets meeting the task requirement serves as a key challenge. Promising future direction includes the development of credible and comprehensive datasets, generalized solutions, explainable claim detection, time-aware claim detection, and prompt-based fine-tuning using large language models.